a body moving with a consonant velocity along a straight line PQR takes 30s to go from P to Q and 10s to go from P to R. If PR is 4m find PQ
1
Expert's answer
2019-11-04T10:11:45-0500

We have:


"v_1=v_2\to \frac{PQ}{t_1}= \frac{PR}{T}"

"\frac{PQ}{30}= \frac{4}{30+10}"

"PQ=3\ m"